OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f * g)(-5)

OpenStudy (anonymous):

A. 23 B. -53 C. 3 D. -9

OpenStudy (zzr0ck3r):

f(-5) = -2(-5)-7 = 10-7 = 3 g(-5) = -4(-5)+3 = 23 23*3 = 69

OpenStudy (zzr0ck3r):

none of the answers are right

OpenStudy (anonymous):

@zzr0ck3r One of those is right!

OpenStudy (zzr0ck3r):

ahh \(f\cdot g \ne f*g\)

OpenStudy (anonymous):

What is the difference?

OpenStudy (zzr0ck3r):

well \((f*g)(x)=f(x)*g(x)\) and \((f\cdot g)(x) = f(g(x))\)

OpenStudy (zzr0ck3r):

we want the latter

OpenStudy (zzr0ck3r):

\(f(g(-5)) = f(23) = -2*23-7=-53\)

OpenStudy (anonymous):

Thanks so much

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

what is f(-7) =?

OpenStudy (anonymous):

No idea.. sorry

OpenStudy (anonymous):

(x^2+6)(-7) ??

OpenStudy (zzr0ck3r):

f(x) = -2x-7 put a -7 where you see an x